# Perceptogram: Reconstructing Visual Percepts from EEG

### v2

@@ Abstract
Visual neural decoding from EEG has improved significantly due to diffusion models that can reconstruct high-quality images from decoded latents. While recent works have focused on relatively complex architectures to achieve good reconstruction performance from EEG, less attention has been paid to the source of this information. In this work, we attempt to discover EEG features that represent perceptual and semantic visual categories, using a simple pipeline. Notably, the high temporal resolution of EEG allows us to go beyond static semantic maps as obtained from fMRI. We show (a) Training a simple linear decoder from EEG to CLIP latent space, followed by a frozen pre-trained diffusion model, is sufficient to decode images with state-of-the-art reconstruction performance. (b) Mapping the decoded latents back to EEG using a linear encoder isolates CLIP-relevant EEG spatiotemporal features. (c) By using other latent spaces representing lower-level image features, we obtain similar time-courses of texture/hue-related information. We thus use our framework, Perceptogram, to probe EEG signals at various levels of the visual information hierarchy. We make our code publicly available: https://github.com/desa-lab/Perceptogram
@@ 1 Introduction
The field of brain decoding from EEG signals has advanced at an extraordinary pace, as methods developed for fMRI have been increasingly applied to EEG. Traditionally, the scope of EEG decoding research has generally been limited to coarse-grained brain state decoding related to tasks (e.g. motor imagery), surprisals (e.g. P300), attention, etc. (Saeidi et al.,, 2021). The direct retrieval of mental content has been understudied due to low signal-to-noise ratio (SNR) (Sadiya et al.,, 2021) and spatial resolution constraints (Burle et al., (2015)).
On the other hand, the machine learning community has focused on complex model designs (e.g transformer-based architecture) and performance scores for visual reconstruction, with less focus on the underlying neural processes Spampinato et al., (2016); Li et al., 2024a . We argue that establishing common ground between these perspectives is essential to advancing the field.
@@ 1.1 Related Work
Basic scientific studies about visual (Kay et al.,, 2008) and semantic representations (Mitchell et al.,, 2008) in the brain with fMRI laid the groundwork for impressive visual reconstructions seen recently (Takagi and Nishimoto,, 2023). Earliest EEG reconstructions suffer from poor experiment design (Li et al.,, 2020) which, combined with the poor quality control in the data, led to inflated reconstruction results due to classifying noise statistics (Spampinato et al.,, 2016). The THINGS-EEG2 dataset (Gifford et al.,, 2022), with improvements in experiment design and quality control, prompted a resurgence of EEG visual decoding Song et al., (2024); Li et al., 2024b . These subsequent studies developed more complicated decoding pipelines, while scientific questions about the organizing principles of the underlying EEG features remained unanswered.
@@
1.2 Motivation
The advent of vision-language foundation models such as CLIP, trained on very large (400M) datasets, has produced rich latent space representations that capture the high-level semantic structure between images Radford et al., (2021). If CLIP embeddings and EEG representations share high-level relational structure, a simple linear mapping between these spaces can be learnt to enable high-performance image reconstruction from the CLIP latent space. This is motivated by the success in fMRI-based visual decoding from Ozcelik and VanRullen, (2023), which forms the basis of our first reconstruction pipeline.
More importantly, we wish to investigate the neurological relevance of EEG features for visual reconstruction. To do this, we learn a linear map from CLIP space to EEG during training, and map decoded CLIP latents back to EEG during evaluation. Intuitively, this decoding-encoding loop, wherein EEG is decoded to latent space and then encoded back, acts as a filter to isolate EEG features (electrodes and time points) carrying shared semantic information. We thus obtain latent-filtered spatio-temporal EEG patterns, or EEG patterns for brevity (in analogy to common spatial patterns from Blankertz et al., (2008) used in Brain-Computer Interfaces). In addition to high-level concepts, our visual experience incorporates lower-level features such as color and texture. What spatio-temporal components of EEG capture these image features? To answer this question, we repeat our analysis by replacing CLIP with other latent spaces that correspond to lower-level visual features and obtain corresponding EEG patterns.
To validate the EEG patterns obtained, we use our pipeline on fMRI data from a different dataset but similar experimental paradigm, to obtain fMRI patterns for various visual features. We find that these patterns from fMRI and EEG are spatially aligned.
@@ 2.1 Dataset
We used the publicly available THINGS-EEG2 (Gifford et al.,, 2022) and Natural Scenes Dataset (NSD) (Allen et al., 2022a, ) for EEG and fMRI analyses, respectively, to validate findings from our EEG analysis. Both datasets are described in the Appendix A.1 and briefly introduced below.
@@ 2.1.2 NSD
7T fMRI data was collected for Microsoft’s COCO images database (Lin et al.,, 2014) with images presented for 3 seconds and 1 second inter-trial interval. 982 images shared across 4 trial-complete subjects were used as the test set, while each subject had 8859 exclusive images (not shared across subjects) that were used as the training set. Image presentation order was pseudo-randomized across the entire image set, with each image presented 3 times to enhance the signal-to-noise ratio.
@@ 2.2.1 Encoder, Decoder
We employ linear regression in both the encoder and decoder, shown to be an effective way to decode latents from fMRI (Ozcelik and VanRullen,, 2023).
@@ 2.2.2 Latent Space
While Fig. 1 illustrates the pipeline using the CLIP latent space, we also use other latent spaces which emphasize different visual features in their reconstructions. With CLIP latents, the reconstructions preserve high-level semantic categories of the images. Latents from VDVAE, PCA or ICA might emphasize other lower-level visual features. While ICA reconstructions emphasize color saturation and contrast, PCA reconstructions capture overall brightness well.
2.2.3 Reconstructer
Reconstructions from CLIP latents require a diffusion module. We have used both Versatile Diffusion, as in Ozcelik and VanRullen, (2023), and unCLIP (Ramesh et al.,, 2022). Images are reconstructed from VDVAE latents using its pretrained frozen decoder, and PCA/ICA simply use linear inverse projections.
@@ 2.3.1 Image reconstruction
We started by using the Versatile Diffusion method of Ozcelik and VanRullen, (2023) but later developed a simpler pipeline using unCLIP (Ramesh et al.,, 2022) which we use for most of our qualitative analyses as it simplifies the overall architecture while maintaining comparable reconstruction performance. A flowchart illustrating image reconstruction for the unCLIP variant of our pipeline is shown in Fig. 2. We introduce an initial VAE encoder into the standard unCLIP framework so our unCLIP diffusion process uses two types of latents as input, an initial (VAE) latent and a conditioning/guidance vector (CLIP).
@@ 2.3.2 Electrode mirroring
To understand how EEG spatially encodes visual features, we perform electrode mirroring experiments. We train the model using unaltered EEG as described in the previous section. During evaluation, we examine reconstructions produced by feeding EEG data mirrored along the midline. For example, during test time, we would swap the EEG between channels O1 and O2 (see Fig. 9 for the electrode topography). We consider two mirrored conditions: In one, we swap the EEG of all non-midline electrodes, and in the other, we swap only EEG for O1 and O2.
@@ 2.3.4 EEG Patterns
This section describes how we obtain the spatio-temporal EEG patterns specific to a visual feature, with textured vs smooth patterns as an example. First, we choose a latent space emphasizing the feature of interest, by manually inspecting reconstructions from various latent spaces. Textures appear to be emphasized in reconstructions from VDVAE, as seen in Fig 5. The encoder and decoder are then trained as described previously. During testing, the decoder is used to first predict image latents from the held-out EEG data, and the encoder is used to project these latents back to EEG space, thus ‘filtering’ the EEG through the chosen latent space. This is illustrated in Fig. 3, which highlights the decoding-encoding loop during testing. Next, we order the reconstructed images along the visual feature of interest (eg. textured vs smooth images). This is done using heuristics described in the Appendix. EEG patterns for all images in each group are averaged, to form the corresponding group EEG pattern (eg. texture pattern vs smooth pattern). Using this procedure, we produce maps for the following visual features: visual semantics (animal vs food vs other), texture (textured vs smooth), hue (red vs blue) and brightness (bright vs dark). In the case of semantic EEG patterns, to minimize color-related confounds, we first grayscaled all images before extracting CLIP latents from them (see Fig. 25).
@@ 2.3.5 EEG-fMRI pattern validation
We repeat the analysis with the NSD dataset to create corresponding fMRI patterns. We use the fMRI patterns as spatial references for the observed EEG patterns, from which we are able to visualize the correlates of EEG activations in the brain source space for the same image presentations.
2.3.6 Qualitative Representational Similarity Analysis (RSA)
To compare the information encoded in different representations (EEG, CLIP, etc), we use representational similarity matrices (RSMs; Kriegeskorte et al., (2008)) generated using the Pearson correlation coefficient between EEG and EEG Patterns corresponding to different stimulus features (e.g., semantic class, texture, luminance, etc.).
2.4 Evaluation
We evaluate image reconstruction performance using the metrics from Ozcelik and VanRullen, (2023) to facilitate direct comparison with state-of-the-art, which are commonly used across many studies. Details are provided in the Appendix.

### v3

# Perceptogram: Reconstructing Visual Percepts and Presumptive Electrode Preference from EEG

@@ Abstract
Visual neural decoding from EEG has improved significantly due to diffusion models that can reconstruct high-quality images from decoded latents. While recent works have focused on relatively complex architectures to achieve good reconstruction performance from EEG, less attention has been paid to the source of this information. We present a unified framework that not only enables image reconstruction from EEG using a simple linear decoder, but also isolates interpretable EEG feature maps that support visual perception. Unlike prior approaches that rely on deep, opaque models, our method leverages the inherent structure of CLIP embeddings to keep the mapping linear. We show that training a simple linear decoder from EEG to CLIP latent space, followed by a frozen pre-trained diffusion model, is sufficient to decode images with state-of-the-art reconstruction performance. Beyond reconstruction, Perceptogram enables the visualization of presumptive electrode preference and EEG patterns, revealing interpretable EEG feature maps that correspond to distinct visual attributes, such as semantic class, texture, and hue. We thus use our framework, Perceptogram, to probe EEG signals at various levels of the visual information hierarchy. We make our code publicly available: https://github.com/desa-lab/Perceptogram
@@ 1 Introduction
Scalp EEG’s low signal-to-noise ratio low_snr and spatial resolution Burle2015 has long constrained its application to basic brain-state decoding, such as motor imagery, attention, and surprisal responses (e.g. P300). While recent advances have extended EEG’s utility to visual decoding, these approaches predominantly rely on deep, non-linear architectures spampinato2016; Li_Wei_Li_Zou_Liu_2024 that, while effective, often obscure the underlying neural representations. This reliance on multi-layered networks has solidified the belief that meaningful visual decoding from EEG demands complex non-linear transformations, prioritizing performance at the cost of interpretability.
We challenge this assumption with Perceptogram – a linear framework that redefines EEG-based visual reconstruction by leveraging the inherent structure of CLIP embeddings through a linear mapping from EEG signals directly into CLIP’s latent space. These latents are then sent to a frozen, pre-trained diffusion model, yielding high-fidelity visual reconstructions. By eliminating the need for deep non-linear stacks, Perceptogram reveals that CLIP’s latent space is intrinsically compatible with the representational dynamics of scalp EEG, enabling accurate and interpretable decoding.
@@ 1.1 Related Work
fMRI studies of visual (Kay_Naselaris_Prenger_Gallant_2008) and semantic representations (mitchell_predicting_2008) laid the groundwork for recent impressive visual reconstructions (takagi_high-resolution_2023). Earlier EEG reconstructions suffered from poor experiment design (Li_Johansen_Ahmed_Ilyevsky_Wilbur_Bharadwaj_Siskind_2020) which, combined with inadequate quality control in the data, led to inflated reconstruction results due to classifying noise statistics (spampinato2016). The THINGS-EEG2 dataset (giffordLargeRichEEG2022), with improvements in experiment design and quality control, prompted a resurgence of EEG visual decoding nice_eeg2024; li2024. These subsequent studies developed more complicated decoding pipelines but did not address the organizing principles of the underlying EEG features.
The advent of vision-language foundation models such as CLIP, trained on very large (400M images) datasets, has produced rich latent space representations that capture the high-level semantic structure between images clipvision. These models enable more efficient encoding of visual-semantic relationships, which has been recently exploited in fMRI-based decoding Ozcelik_VanRullen_2023. Motivated by these findings, we hypothesize that if CLIP embeddings and EEG representations share high-level relational structure, then a simple linear mapping could bridge these spaces for high-performance image reconstruction–circumventing the need for deep, multi-layered networks. This premise forms the foundation of our reconstruction pipeline.
@@
Our Approach
Our approach goes beyond mere image reconstruction from EEG–it serves as a tool to isolate and interpret the EEG features that support the decoding. We do this in two ways: 1. Test-time Perturbation: we systematically perturb EEG inputs (mirroring electrodes, swapping time segments) during test time, and inspect how this changes the original reconstructions. 2. Decoding-Encoding Loop: At the heart of our method is a linear mapping from EEG to CLIP’s latent space. Once decoded, the predicted latents are projected back to EEG space using a linear encoder. Intuitively, this decoding-encoding loop, wherein EEG is decoded to latent space and then encoded back, acts as a filter to isolate EEG features (electrodes and time points) carrying shared semantic information. The result is what we term latent-filtered (spatiotemporal) EEG patterns, directly analogous to common spatial patterns (Blankertz2008) in the Brain-Computer Interface (BCI) literature. To explore the full spectrum of visual processing, we extend this analysis beyond high-level concepts to low-level visual features like color and texture. By replacing CLIP with latent spaces emphasizing these features, we isolate the corresponding spatiotemporal EEG components, uncovering how the brain represents these lower-level visual features. We extend our approach with cross-modal validation using fMRI data from a similar experimental paradigm allowing us to verify the spatial localization of latent-filtered EEG patterns. Together, these components form a unified pipeline that enables both reconstruction and neuroscientific insight into the brain’s dynamic representation of visual categories.
@@ 2.1 Dataset
We used the publicly available THINGS-EEG2 (Gifford_Dwivedi_Roig_Cichy_2022) and Natural Scenes Dataset (NSD) (allen2022massive) for EEG and fMRI analyses, respectively, to validate findings from our EEG analysis. Both datasets are described in the Appendix A.1 and briefly introduced below.
@@ 2.1.2 NSD
7T fMRI data was collected for Microsoft’s COCO images database (coco_dataset_2014) with images presented for 3 seconds and 1 second inter-trial interval. 982 images shared across 4 trial-complete subjects were used as the test set, while each subject had 8859 exclusive images (not shared across subjects) that were used as the training set. Image presentation order was pseudo-randomized across the entire image set, with each image presented 3 times to enhance the signal-to-noise ratio. The original dataset includes the preprocessed betas commonly used in decoding, which is what we used.
@@ 2.2.1 Encoder, Decoder
We employ linear regression in both the encoder and decoder, shown to be an effective way to decode latents from fMRI (Ozcelik_VanRullen_2023). Regularization strength ($\alpha$) is chosen based on the correlation of Subject-1’s predicted latent and the ground truth latent (Table. 1 has all $\alpha$’s).
Table 1: Regularization strengths ($\alpha$) for each latent representation, modality, and mapping direction.
Latent Mapping (modality–direction) $\alpha$
unCLIP CLIP EEG decode 0
unCLIP CLIP EEG encode 1 000
unCLIP CLIP Grayscale EEG decode 1 000
unCLIP CLIP Grayscale EEG encode 1 000
unCLIP VAE EEG decode 0
PCA EEG decode 1 000
PCA EEG encode 1 000 000
ICA EEG decode 1 000
ICA EEG encode 1 000 000
VDVAE EEG decode 1 000
VDVAE EEG encode 1 000
Versatile Diffusion CLIP-Vision EEG decode 1 000
Versatile Diffusion CLIP-Text EEG decode 10 000
unCLIP VAE fMRI decode 100 000
unCLIP CLIP fMRI decode 100 000
unCLIP CLIP fMRI encode 10 000
PCA fMRI decode 100 000
PCA fMRI encode 100 000
ICA fMRI decode 100 000
ICA fMRI encode 100 000
VDVAE fMRI decode 100 000
VDVAE fMRI encode 1 000 000
Versatile Diffusion CLIP-Vision fMRI decode 50 000
Versatile Diffusion CLIP-Text fMRI decode 100 000
unCLIP CLIP MEG decode 10 000
unCLIP VAE MEG decode 10 000
@@ 2.2.2 Latent Space
While Fig. 1 illustrates the pipeline using the CLIP latent space, we also use other latent spaces which emphasize different visual features in their reconstructions. With CLIP latents, the reconstructions preserve high-level semantic categories of the images. Latents from VDVAE, PCA or ICA might emphasize other lower-level visual features.
2.2.3 Reconstructor
Reconstructions from CLIP latents require a diffusion module. We have used both Versatile Diffusion, as in Ozcelik_VanRullen_2023, and unCLIP (ramesh_hierarchical_2022). Images are reconstructed from VDVAE latents using its pretrained frozen decoder, and PCA/ICA simply use linear inverse projections.
2.2.4 Implementation details
Linear models were trained on an AMD EPYC 7302 16-Core Processor using the scikit-learn framework, with each training run completing in a few seconds. Reconstructions using pretrained latent diffusion models were executed on a single NVIDIA A6000 GPU, with an average runtime of approximately 1.5 seconds per image.
2.3 Evaluation
2.3.1 Reconstruction
We used the same performance metrics (see Table 2) as in Ozcelik_VanRullen_2023, which has been used in other followup studies such as MindEye Scotti_Banerjee_Goode_Shabalin_Nguyen_Cohen_Dempster_Verlinde_Yundler_Weisberg_et_al_2023. The 8 metrics we used are Pixel Correlation (PixCorr), Structural Similarity (SSIM), AlexNet layer 2 and 5 outputs pairwise correlations, InceptionNet output pairwise correlation, CLIP ViT output pairwise correlation, EfficientNet output distance, and SwAV output distance. PixCorr and SSIM involve comparing the reconstructed image with the ground-truth (GT) test image. PixCorr is a low-level (pixel) measure that involves vectorizing the reconstructed and GT images and computing the correlation coefficient between the resulting vectors. SSIM is a measure developed by Wang et al. 2004 that computes a match value between GT and reconstructed images as a function of overall luminance match, overall contrast match, and a “structural” match which is defined by normalizing each image by its mean and standard deviation.
2.3.2 Retrieval
To assess the alignment between projected EEG representations and CLIP embeddings, we employed a cross-modal retrieval paradigm. Specifically, we measured Recall@1 and Recall@5, standard retrieval metrics quantifying how often the ground-truth target embedding appears among the top-k nearest neighbors in the retrieval set. For each projected EEG sample, we computed l-2 normalized cosine similarity with all CLIP embeddings in the test set and ranked them by similarity score. Recall@1 reflects the proportion of trials where the correct CLIP embedding was retrieved as the most similar item, while Recall@5 captures retrieval within the top five matches.
@@ 2.4.1 Image reconstruction
We started by using the Versatile Diffusion method of Ozcelik_VanRullen_2023 but later developed a simpler pipeline using unCLIP (ramesh_hierarchical_2022) which we use for most of our qualitative analyses as it simplifies the overall architecture while maintaining comparable reconstruction performance. A flowchart illustrating image reconstruction for the unCLIP variant of our pipeline is shown in Fig. 2. We introduce an initial VAE encoder into the standard unCLIP framework so our unCLIP diffusion process uses two types of latents as input, an initial (VAE) latent and a conditioning/guidance vector (CLIP).
@@
2.4.2 Electrode preference visualization by visually reconstructing the encoding weights
To investigate the semantic tuning properties in our neural-to-CLIP mapping, we performed a detailed analysis of the encoding weights that transform CLIP latent representations into predicted EEG waveforms. These weights provide insight into what visual features maximally drive neural responses at each electrode and time point.
We extracted individual weight vectors from the CLIP-to-EEG encoding matrix, where each vector corresponds to a specific electrode at a given time bin. Each weight vector has the same dimensionality as CLIP embeddings (1024 dimensions) and represents the linear transformation from CLIP space to predicted EEG voltage. To visualize the preferred stimulus features encoded by these weights, we treated each weight vector as if it were a CLIP embedding of an actual image and reconstruct it using unCLIP. This approach reveals the "ideal stimulus" that would maximally activate each electrode according to the learned linear model.
@@ 2.4.3 Electrode mirroring
Motivated by the well known contralateral organization of visual cortex – where stimuli in the right visual field are initially processed by the left visual cortex, and those in the left field by the right cortex – we investigated if we map our EEG to the homologous electrodes across the midline, whether we could see the reconstructed image mirror about the vertical midline. We train the model using unaltered EEG as described in the previous section. During evaluation, we examine reconstructions produced by feeding EEG data mirrored along the midline. For example, during test time, we would swap the EEG between channels O1 and O2 (see Fig. 11 for the electrode topography). We consider two mirrored conditions: In one, we swap the EEG of all non-midline electrodes, and in the other, we swap only EEG for O1 and O2 (over primary visual cortex).
@@ 2.4.5 Latent-filtered EEG Patterns
This section describes how we obtain the spatio-temporal EEG patterns specific to a visual feature, with textured vs smooth patterns as an example. First, we choose a latent space emphasizing the feature of interest, by manually inspecting reconstructions from various latent spaces. Textures appear to be emphasized in reconstructions from VDVAE, as seen in Fig 5. The encoder and decoder are then trained as described previously. During testing, the decoder is used to first predict image latents from the held-out EEG data, and the encoder is used to project these latents back to EEG space, thus ‘filtering’ the EEG through the chosen latent space. This is illustrated in Fig. 3, which highlights the decoding-encoding loop during testing.
The decoder-encoder loop projects the EEG onto the cross-covariance subspace shared between the EEG and the chosen latent space, so only the EEG features covarying with the chosen latent space would be preserved. In an extreme case if the latent space is uncorrelated with the EEG, then the projected EEG will reveal no structure:
Setup.
Let $X\in\mathbb{R}^{n\times d}$ denote EEG features (rows = $n$ classes, columns = $d$ dimensions) and $Y\in\mathbb{R}^{n\times k}$ denote latent embeddings (columns = $k$ dimensions). Define ridge-regularized decoder and encoder
$W_{d}:=(X^{\top}X+\lambda_{d}I_{d})^{-1}X^{\top}Y,\qquad W_{e}:=(Y^{\top}Y+\lambda_{e}I_{k})^{-1}Y^{\top}X.$
Let the composite filter be
$F:=W_{e}^{\top}W_{d}^{\top}.$
Introduce the shorthand
$\Sigma_{xx}:=X^{\top}X+\lambda_{d}I_{d},\quad\Sigma_{yy}:=Y^{\top}Y+\lambda_{e}I_{k},\quad\Sigma_{xy}:=X^{\top}Y,\quad\Sigma_{yx}:=Y^{\top}X=\Sigma_{xy}^{\top}.$ (1)
Lemma 1 (Column space of the composite filter).
With the definitions above,
$\operatorname{col}(F)=\operatorname{col}(\Sigma_{xy}).$
Proof.
First note
$F$ $=W_{e}^{\top}W_{d}^{\top}=(\Sigma_{yy}^{-1}\Sigma_{yx})^{\top}(\Sigma_{xx}^{-1}\Sigma_{xy})^{\top}$
$=\Sigma_{xy}\,\Sigma_{yy}^{-1}\,\Sigma_{yx}\,\Sigma_{xx}^{-1}=\Sigma_{xy}\bigl(\Sigma_{yy}^{-1}\Sigma_{yx}\Sigma_{xx}^{-1}\bigr)=\Sigma_{xy}\,C,$ (2)
where
$C:=\Sigma_{yy}^{-1}\Sigma_{yx}\Sigma_{xx}^{-1}.$ (3)
By the basic fact $\operatorname{col}(AB)\subseteq\operatorname{col}(A)$ for conformable $A,B$, we get
$\operatorname{col}(F)=\operatorname{col}(\Sigma_{xy}C)\subseteq\operatorname{col}(\Sigma_{xy}).$ (4)
Using $\Sigma_{yx}=\Sigma_{xy}^{\top}$,
$F=\Sigma_{xy}\Sigma_{yy}^{-1}\Sigma_{xy}^{\top}\Sigma_{xx}^{-1}.$ (5)
Hence
$\operatorname{rank}(F)=\operatorname{rank}\!\bigl(\Sigma_{xy}\Sigma_{xy}^{\top}\bigr)=\operatorname{rank}(\Sigma_{xy})=:r,$ (6)
since $\operatorname{rank}(AA^{\top})=\operatorname{rank}(A)$ for any AA. By (4) the subspace $\operatorname{col}(F)$ is contained in $\operatorname{col}(\Sigma_{xy})$, and (6) shows they have the same dimension $r$. Therefore $\operatorname{col}(F)=\operatorname{col}(\Sigma_{xy})$. ∎
Next, we order the reconstructed images along the visual feature of interest (eg. textured vs smooth images). This is done using heuristics described in the Appendix. EEG patterns for all images in each group are averaged, to form the corresponding group EEG pattern (eg. texture pattern vs smooth pattern). Using this procedure, we produce maps for the following visual features: visual semantics (animal vs food vs other), texture (textured vs smooth), hue (red vs blue) and brightness (bright vs dark). In the case of semantic EEG patterns, to minimize color-related confounds, we first grayscaled all images before extracting CLIP latents from them (see Fig. 28).
@@ 2.4.6 EEG-fMRI pattern validation
We repeat the analysis with the NSD dataset to create corresponding fMRI patterns which we compare spatially to the latent-filtered EEG patterns.
Representational Similarity Analysis (RSA)
To compare the information encoded in different representations (EEG, CLIP, etc), we use representational similarity matrices (RSMs; kriegeskorteRepresentationalSimilarityAnalysis2008) generated using the Pearson correlation coefficient between EEG and EEG Patterns corresponding to different stimulus features (e.g., semantic class, texture, luminance, etc.).